### v1

@@ 1 Introduction
where $H(\cdot)$ is a convex function defined on a closed convex set $\mathcal{U}\subset\mathbb{R}^{d}$, and $c:\mathcal{X}\rightarrow\mathcal{U}$ is an invertible map (with its inverse denoted by $c^{-1}(\cdot)$). This property is often referred to as hidden convexity and frequently appears in various modern applications, for example, policy optimization in convex reinforcement learning and optimal control [92, 80], generative models [50], supply chain and revenue management [36, 15], training neural networks [83, 30]. In the optimization literature, hidden convexity has been identified much earlier and dates back to at least 1990s in the context of quadratic optimization [78, 8]. Since then, several works have developed various tools to identify such property [7, 6]. More recently, hidden convexity has been established for a wider classes of non-convex programs [81, 84, 14] and for non-monotone games [82, 62], to name just a few. Despite the existence of the convex reformulation, the transformation function $c(\cdot)$ is usually hard to compute or even unknown, and one cannot readily solve the convex reformulation. This motivates the use of (sub-) gradient methods that optimize $F(\cdot)$ directly over the variable $x\in\mathcal{X}$. Perhaps, the most basic algorithm is the projected stochastic (sub-) gradient method (SM). Starting from $x^{0}\in\mathcal{X}$, SM generates a sequence $x^{t}$ via
@@ 1.1 Related work
The most closely related works are [92, 93, 4, 16, 40], which analyze gradient methods under similar structural assumptions in the context of specific applications.
@@
Related structural assumptions. We mention that several other non-convex structural assumptions have been explored in optimization literature that also ensures global convergence, including essential strong convexity [56], quasar (strong) convexity [42], restricted secant inequality [90], error bounds [59], quadratic growth [10], Polyak-Łojasiewicz (PŁ) condition [71, 58] – also known as global Kurdyka-Łojasiweicz (KŁ) and gradient domination condition. For an in-depth discussion on the relationships between these properties, refer to  [47, 72] and references therein. Notably, the PŁ condition along with its various generalizations to constrained minimization such as Proximal-PŁ [47] and variational gradient dominance [85] have gained popularity in the recent years. The convergence of gradient-based methods under the PŁ-type conditions has been extensively analyzed, e.g., in the deterministic setting [47, 89] and the stochastic setting [32, 37, 48, 75, 54, 17, 88, 31]. Despite a few examples [35, 33, 26, 85] that show some variants of the PŁ condition hold, how to verify PŁ-like conditions for non-convex problems remains a big question in general. The situation becomes even more challenging, when dealing with constrained optimization and/or non-differentiable objectives, where a suitable generalization of the gradient dominance needs to be introduced and carefully studied. Unlike the PŁ condition, hidden convexity considered in this work is a very natural property, which easily extends to constrained optimization and non-differentiable objectives.
@@ 2 Hidden Convex Problem Class
Notice that the hidden convex problem class includes the convex problem as a special case when the transformation map $c(\cdot)$ is identical. In addition, it also includes many non-convex problems. For a simple example, let $0<\delta\leq 1$ and consider $\mathcal{X}=[\delta,1]$, $c(x)=x^{2}$, $H(u)=-u$. Then $F(x)=-x^{2}$, albeit concave, is hidden convex on $\mathcal{X}$ by the construction. Another simple example considers $0<\delta<\pi$ and $\mathcal{X}=[\delta,2\pi-\delta]$, $c(x)=cos(x)$, $H(u)=u$. The obtained composition $F(x)=cos(x)$ is also hidden convex on $\mathcal{X}$, although it is both non-convex and non-concave on $\mathcal{X}$.
@@ 2.1 Non-linear least squares [67, 66, 28]
Consider solving a system of nonlinear equations under a box constraint, e.g., $c(x)=0$ with $c(x)=(c_{1}(x),\ldots,c_{d}(x))^{\top}$ for $x\in\mathcal{X}=[0,D]^{d}$, $D>0$. This problem can be equivalently formulated as
$\min_{x\in[0,D]^{d}}\sum_{i=1}^{d}c_{i}^{2}(x)\qquad\text{or }\qquad\min_{x\in[0,D]^{d}}\max_{1\leq i\leq d}|c_{i}(x)|.$
When $c(\cdot)$ is an invertible mapping, it belongs to the hidden convex optimization class. For $d=2$, $c_{1}(x)=x_{1}-1$, and $c_{2}(x)=x_{2}-2x_{1}^{2}+1$, we illustrate its contour plot in fig. 1.
Figure 1: The countour plots of the functions $F(x)=\frac{1}{4}(x_{1}-1)^{2}+\frac{1}{2}(2x_{1}^{2}-x_{2}-1)^{2}$ (top), and $F(x)=\max\left\{\frac{1}{4}|x_{1}-1|,\frac{1}{2}|2x_{1}^{2}-x_{2}-1|\right\}$ (bottom), $x=(x_{1},x_{2})^{\top}$ The left plots present the conour plots in the original space $\mathcal{X}$ and the right plots illustrate the reformulated space $\mathcal{U}$. The red star denotes the global minimum.
@@ 2.2 Minimizing posinomial functions [13, 29]
For power control in communication systems and optimal doping profile problems [13, 29], one often needs to minimize posinomial functions $F(\cdot):\mathbb{R}_{+}^{d}\rightarrow\mathbb{R}$ of the following form
@@
where $H(\cdot)$ is convex. One can easily see that the above problem is hidden convex on a convex compact constraint $\mathcal{X}$.
@@ 2.3 System level synthesis in optimal control [2]
where $\Phi_{\mathbf{x}},\Phi_{\mathbf{u}}\in\mathbb{R}^{(T+1)\times(T+1)}$ are the new variables, $H(\cdot)$ is a strongly-convex function of $\Phi:=(\Phi_{\mathbf{x}},\Phi_{\mathbf{u}})$. $\mathbf{M}\in\mathbb{R}^{(T+1)\times(T+1)}$ is a deterministic matrix, which depends on matrices $A_{t}$, $B_{t}$, $t=0,\ldots,T-1$, and $I$ is the identity matrix. Moreover, there exists a bijection between variables $\mathbf{K}$ and $\Phi$ subject to the constraints of the reformulated problem. The (inverse of the) map $c(\cdot)$ is given by $\mathbf{K}=c^{-1}(\Phi):=\Phi_{\mathbf{u}}\Phi_{\mathbf{x}}^{-1}$ [2]. Therefore, one can easily verify that the optimization problem over $\mathbf{K}$ is hidden convex. A number of other problems in optimal control also admit suitable convex reformulations. We refer readers to [12, 80] for more examples.
@@ 2.4 Convex reinforcement learning [92]
Notice that $F(\cdot)$ is not convex in $\pi$ in general. However, for several commonly used utility functions, $H(\cdot)$ exhibits convexity in the occupancy measure $\lambda^{\pi}$. For standard RL, $H(\lambda^{\pi})=r^{\top}\lambda^{\pi}$ is linear in $\lambda^{\pi}$, where $r$ is the reward vector. For the pure exploration setting, focused on fully exploring the transitions in the environment, $H(\lambda^{\pi})$ represents the negative entropy of $\lambda^{\pi}$, which is also convex  [92]. For the imitation learning where the objective is to imitate the expert’s behavior given their sampled trajectories, $H(\lambda^{\pi})$ denotes the KL-divergence between $\lambda^{\pi}$ and the state-action occupancy measure learned from the expert’s sampled trajectories, which is also convex [92]. Thus, the convex RL problem belongs to the hidden convex class with $\mathcal{X}=\Pi$ and $c(x)=\lambda^{\pi}$ (with $x=\pi$). Under mild assumptions on the initial distribution $\rho$, the constant $\mu_{c}>0$ can be estimated, see e.g., [92]. Note that in convex RL, we can control $\lambda^{\pi}$ only implicitly by changing the policy $\pi$. The exact computation of the transformation map and its inverse requires the knowledge of the state transition probability kernel and can be computationally expensive.
@@ 2.5 Revenue management and inventory control [15, 16]
where $d$ denotes the number of demand classes in the airline networks, $x\in\mathbb{R}^{d}$ is the booking limit control threshold for each demand class, $\xi$ is the random demand vector (of the same dimension as $x$) during the reservation stage, $x\wedge\xi$ denotes the number of reservations accepted, and $r^{\top}(x\wedge\xi)$ denotes revenue collected during the reservation stage with $r\in\mathbb{R}^{d}$ being the price vector. In the service stage, $\Gamma(x\wedge\xi,\eta)$ denotes the penalty on the airline companies when there are $x\wedge\xi$ number of reservations with plane seats capacity $\eta$ that is random, $w$ is the actual number of passengers that can get on the plane, $l$ is the penalty vector for declining passengers with reservation to get on the plane. Notice that $F$ is non-convex in $x$ due to the truncation between $x$ and $\xi$. However, when $\xi$ admits component-wise independent coordinates, this problem admits a convex reformulation via a variable change [15], i.e., $u=c(x)=\mathbb{E}_{\xi}[x\wedge\xi]$. Note that comparing to previous applications, the transformation function involves unknown distribution and thus is not explicitly known.
@@ 3.1 Globally optimal solution
Let $F(\cdot)$ be hidden convex on $\mathcal{X}$ and $\bar{x}\in\mathcal{X}$ be a stationary point. If the map $c(\cdot)$ is differentiable at $\bar{x}$ , then $\bar{x}$ is a global minimum, i.e., $F(\bar{x})\leq F(x)$ for any $x\in\mathcal{X}$.
@@
where $\bar{u}=c(\bar{x})$. As the map $c(\cdot)$ is invertible, then $\nabla c(\bar{x})y=0$ for some $y\in\mathbb{R}^{d}$ implies $y=0$. Thus, we have $0\in\partial_{u}H(\bar{u})+\partial_{u}\delta_{\mathcal{U}}(\bar{u})$. Since function $H(\cdot)$ is convex, by the sufficient optimality condition, $\bar{u}$ is a globally optimal solution, i.e., $H(\bar{u})\leq H(u)$ for any $u\in\mathcal{U}$. As a result, we have $F(\bar{x})=H(\bar{u})\leq H(u)=F(x)$ for any $x\in\mathcal{X}$. ∎
Note that a similar result appeared in [92] under additional smoothness assumptions on $H(\cdot)$ and $c(\cdot)$. The above proof is much simpler and does not require smoothness.
@@ 3.2 Connections with gradient dominated functions
It is natural to ask what is the connection between hidden convex problems and previously studied gradient dominated function classes that also ensure the global convergence of gradient-based algorithms in the smooth setting [47]. Unfortunately, the exact characterization is difficult to establish in the constrained setting. With the following proposition, we show that a problem satisfies the global KŁ condition if it is hidden strongly convex ($\mu_{H}>0$).
@@
Let $F(\cdot)$ be differentiable and $(\mu_{c},\mu_{H})$-hidden strongly convex, the map $c(\cdot)$ be differentiable on $\mathcal{X}$. Then the optimization problem satisfies the global KŁ condition, i.e.,
$\min_{h_{x}\in\partial\delta_{\mathcal{X}}(x)}\left\|\nabla F(x)+h_{x}\right\|^{2}\geq 2\mu_{H}\mu_{c}^{2}\left(F(x)-F^{*}\right)\qquad\text{for all }x\in\mathcal{X}.$ (11)
@@
Since $H(\cdot)$ is differentiable and strongly convex on a convex set $\mathcal{U}$, then it satisfies the following Proximal-PŁ condition [47] (Appendix F)
$\mathcal{D_{\delta_{\mathcal{U}}}}(u,\mu_{H})\geq 2\mu_{H}(H(u)-H(u^{*}))\qquad\text{for all }u\in\mathcal{U},$
where $\mathcal{D_{\delta_{\mathcal{U}}}}(u,\mu_{H}):=-2\mu_{H}\min_{v\in\mathcal{U}}\left\{\langle\nabla H(u),v-u\rangle+\frac{\mu_{H}}{2}\left\|u-v\right\|^{2}\right\}$. The inequality above implies global KŁ condition, see Appendix G in [47],
$\min_{h_{u}\in\partial\delta_{\mathcal{U}}(u)}\left\|\nabla H(u)+h_{u}\right\|^{2}\geq 2\mu_{H}\left(H(u)-H(u^{*})\right).$
Combined with the chain rule, we have for any $u=c(x)$
$\min_{h_{x}\in\partial\delta_{\mathcal{X}}(x)}\left\|\nabla F(x)+h_{x}\right\|^{2}$ $=$ $\min_{h_{u}\in\partial\delta_{\mathcal{U}}(u)}\left\|\nabla c(x)(\nabla H(u)+h_{u})\right\|^{2}\geq 2\mu_{H}\mu_{c}^{2}\left(F(x)-F^{*}\right),$
where $F^{*}=F(x^{*})$. This concludes the proof. ∎
Note that even in this restrictive case when $\mu_{H}>0$, it remains elusive how condition (11) can be used to establish global convergence of Projected SGD. To our knowledge, under the condition (11), no global analysis of stochastic gradient methods exists in the literature even for smooth $F(\cdot)$. In the more interesting case when $H(\cdot)$ is merely convex, the above global KŁ condition becomes vacuous as $\mu_{H}=0$. Despite of this, Proposition 1 still holds even when $\mu_{H}=0$.

### v2

@@ 1 Introduction
where $H(\cdot)$ is a convex function defined on a closed convex set $\mathcal{U}\subset\mathbb{R}^{d}$, and $c:\mathcal{X}\rightarrow\mathcal{U}$ is an invertible map (with its inverse denoted by $c^{-1}(\cdot)$). This property is often referred to as hidden convexity and frequently appears in various modern applications, for example, policy optimization in convex reinforcement learning and optimal control [107, 94, 102], generative models [58], supply chain and revenue management [40, 18], training neural networks [97, 34]. In the optimization literature, hidden convexity has been identified much earlier and dates back to at least 1990s in the context of quadratic optimization [92, 11]. Since then, several works have developed various tools to identify such property [10, 9]. More recently, hidden convexity has been established for a wider classes of non-convex programs [95, 98, 17] and for non-monotone games [96, 74], to name just a few. The effect of general smooth parameterizations on the landscape of non-convex problems was recently studied in [62].
Despite the existence of the convex reformulation, the transformation function $c(\cdot)$ is usually hard to compute or even unknown, and one cannot readily solve the convex reformulation. This motivates the use of (sub-) gradient methods that optimize $F(\cdot)$ directly over the variable $x\in\mathcal{X}$. Perhaps, the most basic algorithm is the projected stochastic (sub-) gradient method (SM). Starting from $x^{0}\in\mathcal{X}$, SM generates a sequence $x^{t}$ via
@@ 1.1 Related work
Perhaps, one of the first works using hidden convexity to analyze iterative methods is [79], which established convergence rates of cubic regularized Newton method. However, it remains unclear how to extend their results to non-smooth or stochastic setting. The most closely related to our work are [107, 108, 7, 19, 45], which analyze gradient methods under similar structural assumptions in the context of specific applications.
@@
Related structural assumptions. We mention that several other non-convex structural assumptions have been explored in optimization literature that also ensure global convergence, including essential strong convexity [68], quasar (strong) convexity [48], restricted secant inequality [105], error bounds [71], quadratic growth [13], Polyak-Łojasiewicz (PŁ) condition [85, 70] – also known as global Kurdyka-Łojasiweicz (KŁ) or gradient domination condition. For an in-depth discussion on the relationships between these properties, refer to  [55, 86] and references therein. Notably, the PŁ condition along with its various generalizations to constrained minimization such as Proximal-PŁ [55] and variational gradient dominance [99] have gained popularity in the recent years. The convergence of gradient-based methods under the PŁ-type conditions has been extensively analyzed, e.g., in the deterministic setting [55, 104] and the stochastic setting [36, 41, 56, 89, 65, 20, 103, 35]. Unfortunately, despite a few examples [39, 37, 30, 99] that show some variants of the PŁ condition hold, how to verify PŁ-like conditions for non-convex problems remains a big question in general. The situation becomes even more challenging, when dealing with constrained optimization and/or non-differentiable objectives, where a suitable generalization of the gradient dominance needs to be introduced and carefully studied. For a non-differentiable $F$, convergence results under the gradient domination condition often make very strong algorithmic dependent assumptions (e.g., strong descent property), and are difficult to extend to stochastic setting [1, 5, 6, 42]. In the constrained setting, a suitable generalization of the gradient dominance is often algorithmic dependent. Specifically, one needs to replace the gradient norm by a suitable generalized notion of stationarity, e.g. Frank-Wolfe gap or gradient mapping [99]. However, verifying such algorithmic dependent assumptions can be difficult in specific applications. Unlike the gradient domination condition, the hidden convexity considered in this work is a very natural property, which easily extends to constrained optimization and non-differentiable objectives.
@@ 2 Hidden Convex Problem Class
Remarks on condition C.2. First, we notice that (6) is equivalent to $1/\mu_{c}$-Lipschitz continuity of the inverse map $c^{-1}$, i.e.,
$\left\|c^{-1}(u)-c^{-1}(v)\right\|\leq\mu_{c}^{-1}\left\|u-v\right\|\qquad\text{for all }u,v\in\mathcal{U}.$ (7)
Second, if the inverse map $c^{-1}(\cdot)$ is continuously differentiable [^1] This can be verified using a global inverse function theorem [59, 52] if $c(\cdot)$ is continuously differentiable. on $\mathcal{U}$, then there are simple necessary and sufficient conditions, which can help to verify (6). Using standard arguments (see, e.g., Section 1.1.2. in [83] or Section 1.2.2. in [78]), we can show that (7) is equivalent to
$\left\|\nabla c^{-1}(u)\right\|_{\text{op}}\leq 1/\mu_{c}\qquad\text{for any }u\in\mathcal{U}.$ (8)
When $c(\cdot)$ is continuously differentiable on $\mathcal{X}$, condition (6) implies $\left\|\nabla c(x)z\right\|\geq\mu_{c}\left\|z\right\|$ for any $x\in\mathcal{X}$ and $z\in\mathbb{R}^{d}$, see the proof of Proposition 2 for details. We remark that, in general, the map in item C.2 is not required to be continuously differentiable and the results of Section 4 do not require a differentiable $c(\cdot)$.
Simple examples. Notice that the hidden convex problem class includes the convex problem as a special case when the transformation map $c(\cdot)$ is identical. In addition, it also includes many non-convex problems. For a simple example, let $0<\delta\leq 1$ and consider $\mathcal{X}=[\delta,1]$, $c(x)=x^{2}$, $H(u)=-u$. Then $F(x)=-x^{2}$, albeit concave, is hidden convex on $\mathcal{X}$ by the construction. Another simple example considers $0<\delta<\pi$ and $\mathcal{X}=[\delta,2\pi-\delta]$, $c(x)=\cos(x/2)$, $H(u)=2u^{2}-1$. The obtained composition $F(x)=2\cos^{2}(x/2)-1=\cos(x)$ is hidden strongly convex on $\mathcal{X}$ with $\mu_{c}=\sin(\delta/2)/2$, although it is both non-convex and non-concave on $\mathcal{X}$. From the above toy example, it already becomes clear that the condition $\mu_{c}>0$ in item C.2 is necessary for global convergence of Projected GD (e.g., merely assuming an invertible $c(\cdot)$ is not sufficient). Indeed, by taking $\delta=0$, we have $\mu_{c}=0$, but this generates a spurious stationary point at $x=0$, which is also a maxima, where GD gets stuck.
@@ 2.1 Non-linear least squares [80, 79, 32]
Consider solving a system of nonlinear equations, e.g., $c(x)=0$ with $c(x)=(c_{1}(x),\ldots,c_{d}(x))^{\top}$ for $x\in\mathbb{R}^{d}$. This problem can be equivalently formulated as
$\min_{x\in\mathbb{R}^{d}}\sum_{i=1}^{d}c_{i}^{2}(x)\qquad\text{or }\qquad\min_{x\in\mathbb{R}^{d}}\max_{1\leq i\leq d}|c_{i}(x)|.$ (9)
When $c(\cdot)$ is invertible and condition (6) holds, it belongs to the hidden convex optimization class. Despite having a benign reformulation, such problems can be extremely difficult to solve to global optimum. For example, Jarre [53] study the following function
$F(x)=\frac{1}{4}(x_{1}-1)^{2}+c\sum_{i=1}^{d-1}(x_{i+1}-2x_{i}^{2}-1)^{2}$ (10)
for some constant $c>0$ (independent of dimension). Despite being hidden convex and having a unique minima, it is shown that any descent method initialized at $x^{0}=(-1,1,\ldots,1)$ takes an exponential number of iterations (in dimension $d$) to reduce the function value by $75\%$.
In practice, solving problems of type (9) can be challenging even when the dimension is moderate and other challenges such as non-smoothness of $c(x)$ and $H(u)$ persist, see e.g., non-smooth variants of (10) in [47]. In fig. 1, we illustrate the contour plot of such functions and their convex reformulations for dimension $d=2$.
Figure 1: The contour plots of the functions $F(x)=\frac{1}{4}(x_{1}-1)^{2}+\frac{1}{2}(2x_{1}^{2}-x_{2}-1)^{2}$ (top), and $F(x)=\max\left\{\frac{1}{4}|x_{1}-1|,\frac{1}{2}|2x_{1}^{2}-x_{2}-1|\right\}$ (bottom), $x=(x_{1},x_{2})^{\top}$ The left plots present the contour plots in the original space $\mathcal{X}$ and the right plots illustrate the reformulated space $\mathcal{U}$. The red star denotes the global minimum.
@@ 2.2 Minimizing posinomial functions [16, 33]
For power control in communication systems and optimal doping profile problems [16, 33], one often needs to minimize posinomial functions $F(\cdot):\mathbb{R}_{++}^{d}\rightarrow\mathbb{R}$ of the following form
@@
where $H(\cdot)$ is convex.
We can verify that the above problem is hidden convex on a convex compact set $\mathcal{X}\subset\mathbb{R}_{++}^{d}$. Specifically, notice that for any $i=1,\ldots,d$ and $x\in\mathcal{X}$, the function $\log(x_{i})$ is $(\max_{x\in\mathcal{X}}x_{i})^{-1}$-strongly monotone. Therefore, condition (C.2) holds with $\mu_{c}=(\max_{x\in\mathcal{X}}\left\|x\right\|_{\infty})^{-1}$. [^2] We remark that in this example the parameter $\mu_{c}$ depends on the upper limit of the variables $x$. On the other hand, the lower limit, $\min_{1\leq i\leq d}\min_{x\in\mathcal{X}}x_{i}$, influences the diameter $D_{\mathcal{U}}$ of the set $\mathcal{U}=c(\mathcal{X})$. This quantity will also frequently appear in the complexity guarantees in Sections 4, 5, and 6.
@@ 2.3 System level synthesis in optimal control [3]
where $\Phi_{\mathbf{x}},\Phi_{\mathbf{u}}\in\mathbb{R}^{(T+1)\times(T+1)}$ are the new variables, $H(\cdot)$ is a strongly-convex function of $\Phi:=(\Phi_{\mathbf{x}},\Phi_{\mathbf{u}})$. $\mathbf{M}\in\mathbb{R}^{(T+1)\times(T+1)}$ is a deterministic matrix, which depends on matrices $A_{t}$, $B_{t}$, $t=0,\ldots,T-1$, and $I$ is the identity matrix. Moreover, there exists a bijection between variables $\mathbf{K}$ and $\Phi$ subject to the constraints of the reformulated problem. The (inverse of the) map $c(\cdot)$ is given by $\mathbf{K}=c^{-1}(\Phi):=\Phi_{\mathbf{u}}\Phi_{\mathbf{x}}^{-1}$ [3]. Therefore, we recognize the hidden convex structure (4) with invertible $c(\cdot)$.333 Note that, however, it remains elusive whether the condition (6) holds across the entire domain. It is possible that a sufficient condition such as (8) can be verified on the trajectory of some optimization methods instead. A number of other problems in optimal control also admit suitable convex reformulations. We refer readers to [15, 94] for more examples.
@@ 2.4 Convex reinforcement learning [107]
Notice that $F(\cdot)$ is not convex in $\pi$ in general. However, for several commonly used utility functions, $H(\cdot)$ exhibits convexity in the occupancy measure $\lambda^{\pi}$. For standard RL, $H(\lambda^{\pi})=r^{\top}\lambda^{\pi}$ is linear in $\lambda^{\pi}$, where $r$ is the reward vector. For the pure exploration setting, focused on fully exploring the transitions in the environment, $H(\lambda^{\pi})$ represents the negative entropy of $\lambda^{\pi}$, which is also convex  [107]. For the imitation learning where the objective is to imitate the expert’s behavior given their sampled trajectories, $H(\lambda^{\pi})$ denotes the KL-divergence between $\lambda^{\pi}$ and the state-action occupancy measure learned from the expert’s sampled trajectories, which is also convex [107]. Thus, the convex RL problem belongs to the hidden convex class with $\mathcal{X}=\Pi$ and $c(x)=\lambda^{\pi}$ (with $x=\pi$). Under mild assumptions on the initial distribution $\rho$, the constant $\mu_{c}>0$ can be estimated, see e.g., Proposition H.1. in [107]. Note that in convex RL, we can control $\lambda^{\pi}$ only implicitly by changing the policy $\pi$. The exact computation of the transformation map and its inverse requires the knowledge of the state transition probability kernel and can be computationally expensive.
@@ 2.5 Revenue management and inventory control [18, 19]
where $d$ denotes the number of demand classes in the airline networks, $x\in\mathbb{R}^{d}$ is the booking limit control threshold for each demand class, $\xi$ is the random demand vector (of the same dimension as $x$) during the reservation stage, $x\wedge\xi$ denotes the number of reservations accepted, and $r^{\top}(x\wedge\xi)$ denotes revenue collected during the reservation stage with $r\in\mathbb{R}^{d}$ being the price vector. In the service stage, $\Gamma(x\wedge\xi,\eta)$ denotes the penalty on the airline companies when there are $x\wedge\xi$ number of reservations with plane seats capacity $\eta$ that is random, $w$ is the actual number of passengers that can get on the plane, $l$ is the penalty vector for declining passengers with reservation to get on the plane. Notice that $F$ is non-convex in $x$ due to the truncation between $x$ and $\xi$. When $\xi$ admits component-wise independent coordinates and $x<\text{esssup}\,\xi$, this problem admits a convex reformulation via a variable change [40, 18], i.e., $u=c(x)=\mathbb{E}_{\xi}[x\wedge\xi]$. Note that the upper bound on the booking limit threshold $D<\text{esssup}\,\xi$, the largest possible demands, naturally holds in revenue management and implies that Condition C.2 holds. Note that comparing to previous applications, the transformation function involves unknown distribution and thus is not explicitly known.
@@ 3.1 Globally optimal solution
Let $F(\cdot)$ be weakly convex and hidden convex on $\mathcal{X}$ with $\bar{x}\in\mathcal{X}$ being its stationary point. If the map $c(\cdot)$ is differentiable at $\bar{x}$, then $\bar{x}$ is a global minimum, i.e., $F(\bar{x})\leq F(x)$ for any $x\in\mathcal{X}$.
@@
where $\bar{u}=c(\bar{x})$. As the map $c(\cdot)$ is invertible with a Lipschitz continuous inverse by (C.2), then its Jacobian $\nabla c(x)$ is invertible at $\bar{x}$ (see e.g., Corollary 3.3. in [61]). Therefore, (14) implies that $0\in\partial_{u}H(\bar{u})+\partial_{u}\delta_{\mathcal{U}}(\bar{u})$. Since function $H(\cdot)$ is convex, by the sufficient optimality condition, $\bar{u}$ is a globally optimal solution, i.e., $H(\bar{u})\leq H(u)$ for any $u\in\mathcal{U}$. As a result, we have $F(\bar{x})=H(\bar{u})\leq H(u)=F(x)$ for any $x\in\mathcal{X}$. ∎
Note that a similar result appeared in Theorem 4.2 of [107] under additional smoothness assumption on $c(\cdot)$. The above proof is much simpler and does not require smoothness.
@@ 3.2 Connections with gradient dominated functions
It is natural to ask what is the connection between hidden convex problems we consider and previously studied gradient dominated functions [1, 5, 6, 42] that can also be used to ensure the global convergence of (sub-)gradient methods. We will distinguish between the weak (15) and the strong (16) variants of gradient domination condition. It turns out that hidden convexity implies the weak gradient domination and hidden strong convexity implies its strong version in the sense specified below.
@@
Let $F(\cdot)$ be weakly convex and hidden convex with modulus $\mu_{c}$ on $\mathcal{X}$, the map $c(\cdot)$ be continuously differentiable on $\mathcal{X}$.
$(i)$ If the set $\mathcal{U}=c(\mathcal{X})$ is bounded with diameter $D_{\mathcal{U}}$, then
$\inf_{s_{x}\in\partial(F+\delta_{\mathcal{X}})(x)}\left\|s_{x}\right\|\geq\frac{\mu_{c}}{D_{\mathcal{U}}}\left(F(x)-F^{*}\right)\qquad\text{for all }x\in\mathcal{X}.$ (15)
$(ii)$ If $F(\cdot)$ is $(\mu_{c},\mu_{H})$-hidden strongly convex, then
$\inf_{s_{x}\in\partial(F+\delta_{\mathcal{X}})(x)}\left\|s_{x}\right\|^{2}\geq 2\mu_{H}\mu_{c}^{2}\left(F(x)-F^{*}\right)\qquad\text{for all }x\in\mathcal{X}.$ (16)
@@
$(i)$ Since the map $c(\cdot)$ is invertible with a Lipschitz continuous inverse by (C.2), its Jacobian $\nabla c(x)$ is invertible for all $x\in\mathcal{X}$ (Corollary 3.3. in [61]). Thus, $\left\|(\nabla c(x))^{-1}\right\|_{\text{op}}$ is bounded on a compact set $\mathcal{X}$ and we can apply a global inverse function theorem due to Hadamard, see e.g., Theorem 1 in [52]. It implies that the inverse map $c^{-1}(\cdot)$ is also continuously differentiable on $\mathcal{U}$. Therefore, we can show that condition (8) holds, i.e., $\left\|\nabla c^{-1}(u)\right\|_{\text{op}}\leq 1/\mu_{c}$ for all $u\in\mathcal{U}$. By the sub-gradient inequality applied to $H+\delta_{\mathcal{U}}$, we have for any $x\in\mathcal{X}$, $x^{*}\in\mathcal{X}^{*}$, $u=c(x)$, $u^{*}=c(x^{*})$ and any $s_{u}\in\partial(H+\delta_{\mathcal{U}})(u)$ that
$F(x)-F(x^{*})$ $=$ $H(u)-H(u^{*})$
$\leq$ $\langle s_{u},u-u^{*}\rangle$
$\leq$ $\|s_{u}\|\cdot\|u-u^{*}\|$
$=$ $\|\nabla c^{-1}(u)\,s_{x}\|\cdot\|u-u^{*}\|$
$\leq$ $\|\nabla c^{-1}(u)\|_{\text{op}}\cdot\|s_{x}\|\cdot\|u-u^{*}\|$
$\leq$ $\frac{D_{\mathcal{U}}}{\mu_{c}}\|s_{x}\|,$
where the first inequality uses convexity of $H$, the second equality uses the chain rule with $s_{x}\in\partial(F+\delta_{\mathcal{X}})(x)$. Taking the infimum over $s_{x}$ concludes the proof of the first statement.
$(ii)$ First, we will show that C.2 implies $\left\|\nabla c(x)\,z\right\|\geq\mu_{c}\left\|z\right\|$ for any $z\in\mathbb{R}^{d}$. Let $x\in\operatorname{int}\mathcal{X}$, then there exists $\alpha>0$ such that for any $\tau\in[0,\alpha]$ the point $x+\tau z\in\mathcal{X}$. Thus, we can write
[…]
Using (6), we have
[…]
Dividing by $\alpha>0$ and taking the limit $\alpha\rightarrow 0$, we arrive at $\|\nabla c(x)z\|\geq\mu_{c}\|z\|$ for all $z\in\mathbb{R}^{d}$. If $x\in\partial\mathcal{X}$, then there exists a sequence $\left\{x_{t}\right\}_{t\geq 0}\subset\operatorname{int}\mathcal{X}$ such that $\lim_{t\rightarrow\infty}x_{t}=x$. By continuity of $\|\nabla c(x)z\|$ in $x$, we have $\|\nabla c(x)z\|=\lim_{t\rightarrow\infty}\|\nabla c(x_{t})z\|\geq\mu_{c}\|z\|$.
Next, by strong convexity of $H+\delta_{\mathcal{U}}$, we have for any $s_{u}\in\partial(H+\delta_{\mathcal{U}})(u)$ that
$H(v)\geq H(u)+\langle s_{u},v-u\rangle+\frac{\mu_{H}}{2}\left\|v-u\right\|^{2}\qquad\text{for any }u,v\in\mathcal{U}.$
Minimizing both sides over $v\in\mathcal{U}$, we get
$\left\|s_{u}\right\|^{2}\geq 2\mu_{H}\left(H(u)-H(u^{*})\right).$ (17)
Finally, using the chain rule, we have for any $u=c(x)$
$\left\|s_{x}\right\|^{2}$ $=$ $\left\|\nabla c(x)\,s_{u}\right\|^{2}\geq\mu_{c}\left\|s_{u}\right\|^{2}\geq 2\mu_{H}\mu_{c}^{2}\left(F(x)-F(x^{*})\right),$
where in the last inequalities, we used the fact that $\left\|\nabla c(x)\,z\right\|\geq\mu_{c}\left\|z\right\|$ for any $z\in\mathbb{R}^{d}$, and (17). It only remains to take the infimum over $s_{x}$.
In certain special cases, one can establish convergence rates of gradient methods using such gradient dominance conditions. In particular, if there is no noise and the sequence $\{x^{t}\}_{t\geq 0}$ satisfies a (strong) descent property, the results from [1] indicate global convergence rates. However, given that in general the sub-gradient method does not have a descent property, this implication seems limiting even in the deterministic case. When $F$ is additionally smooth, existing theory of Projected SGD under gradient dominated conditions is limited either to unconstrained case [41, 36] with the strong gradient dominance condition (16) [55, 67], or requires to replace (15) with a weak Proximal-PŁ inequality [22], which is different from (15). Unfortunately, the precise connection between (15) and weak Proximal-PŁ appears difficult to establish and is beyond the scope of the current work.
To circumvent the above mentioned technical difficulties, in what follows, we will not use the result of Proposition 2, but develop a direct convergence proof utilizing hidden convexity. Our proof technique allows to handle non-smooth, stochastic and constrained cases automatically without selecting a proper variant of PŁ condition (based on algorithm) and attempting to verify it.